# The Chemistry and Physics of 199Hg Nuclear Spin Polarization Relaxation in Quantum Magnetometry Cells

### v1

@@ I.5.1 Hydrogen surface states
Again, using the WKB approximation, the surface states of an H atom can be determined. The van der Waals coefficient is 2/3 that of Hg because the principal contributing optical transition for H is at a shorter wavelength. Also, the hard-sphere radius of H is smaller than that of Hg, so we can take $d_{0}\approx 0.12$ nm. Together with its smaller mass, this leads to the result that H has only four bound states at 0.08, 0.03, 0.01, and 0.003 $e$V. Although the energies are smaller than for Hg, the $\epsilon$ is enhanced by a relatively large $\lambda_{T}$ compared to Hg. The net result is $\epsilon=2\times 10^{-6}$, about an order of magnitude smaller than for Hg. Nevertheless, this result implies a microscopically long time for an excited Hg atom to interact with H atoms temporarily stuck to the surface.
We can estimate the effective pressure of H on the surface, assuming that a fraction $62\times 10^{-}$ 6 of H atoms are imprisoned in a region about 0.3 nm thick. If there is 1 mbar H gas in a cell, the number density in the thin region near the cell wall corresponds to a pressure of about 4 bar. Alternatively, with a pressure $P_{0}$ in cell volume, the pressure near the surface (just like the isothermal atmosphere) should be $P_{0}e^{E_{max}/k_{b}T}=P_{s}$, and for $P_{0}=0.001$ bar, with an approximate binding energy of 0.21 $e$ V, it leads to $P_{s}=4.4$ bar and is astonishingly close to the previous estimate of 4 bar.
Relaxation on the wall can be due to an enhanced interaction with H atoms or with the open bonds created when a surface is attacked by atomic H or excited Hg atoms. These open (covalent) bonds carry roughly an electron magnetic moment. The magnetic interaction that leads to relaxation is presented in [38] and can be easily modified for other magnetic moments.
@@ I.6 Relaxation in the Gas phase
Paramagnetic atoms and molecules, which include oxygen and free radicals, can exist in the buffer gas as impurities or the result of 254 nm radiation. Subsequent reactions with excited H can release atomic H (an unpaired spin, so technically a free radical) that exists in the states $F=0$ and $F=1$, the latter that exists as 3/4 of the total H and presents an effective 1/2 of an electron spin due to the addition of the proton nuclear angular momentum.
The relaxation of nuclear spin $I=1/2$ 1S0 atoms by gaseous paramagnetic atoms is theoretically analyzed by Jameson, Jameson, and Hwang [46] for the specific case of O2 and 129Xe, and further developed in [47], for both O2 on 3He and on 129Xe Rewritten here in terms of SI units, and more generally for a colliding molecule (or atom) $X$ with total angular momentum $F$ representing the sum of the internal angular momentum of the constituent nuclei and electron spins of $X$,
@@
$\gamma_{I}$ and $\gamma_{e}$ are the gyromagnetic ratios for the $I=1/2$ nucleus, and for the bare electron respectively; $g_{F}$ is the Landè $g-$ factor for $X$, and $\mu_{0}=4\times 10^{-7}$. $d$ is the characteristic intermolecular interaction length, $\mu$ is the reduced mass of the colliding species, $k_{B}$ is the Boltzmann constant and $T$ is the absolute temperature (the factor in parentheses is the reciprocal of the mean relative velocity). $n_{0}$ is Loshmidt constant ($n_{0}=2.69\times 10^{25}$/m3 corresponding to the STP number density of an ideal gas), and $[X]$ is the density of the gas $X$ in Amagats. The factor $F(U,T)$ is determined by the interaction potential $U$ between the colliding species, and is unity at high temperature in the hard-core limit. In general, it is determined by integrating solutions of the radial Schr"odinger equation with the potential $U$, together with an average over the Boltzmann momentum distribution. It is noted in [47] that the hyperfine Fermi contact term, proportional to $|\psi(0)|^{4}$, should generally be included in $F(V,T)$, where $\psi(0)$ is the amplitude of the electron wave function of the colliding molecule at the nucleus $I$. The interaction time (correlation time) for gas phase molecules is very short, so effects due to an applied magnetic field become small when the electron precession over the collision time results in a small directional change in the electron(s) spin angle(s); therefore, in all cases under consideration such effects can be ignored for the gas phase.
@@
In [47], the experimental relaxation rate for 3He in O2 is reported to be approximately $0.5$ [s Amagat], which implies $F(U,T)=1.7$. For 129Xe, $F(U,T)\approx 6$ is estimated to be [46] which implies a relaxation rate of 0.7/[s Amagat].
@@
Data in [48] indicates that the relaxation rate of 199Hg in O2 is 800/[s Amagat], which about 1000 times faster than for the inert gases 3He and 129Xe. Using the analysis for inert gases implies that $F(U,T)=1.5\times 10^{4}$ although the two valence electrons of 199Hg likely enhance a hyperfine interaction between the atom and the molecule, the analysis for inert gases is not adequate. Another possibility is that van der Waals molecules can be formed, in which case the needed three-body collisions would make the relaxation rate pressure dependent. This would be interesting to study.
@@
We can try to apply this result to the relaxation of 199Hg by H. Because H has a nuclear spin of 1/2, 3/4 of the atoms will be in the hyperfine level with $F=1$ and $\gamma_{s}$ is reduced by a factor of two compared to O2. Comparing $F(U,T)$ with previous cases of inert gases suggests that $F(U,T)$ scales as the square root of the reduced mass, which we might expect from the Schrödinger equation. Scaling by $\mu$ and $\gamma_{s}^{2}$, and by the fraction of atoms in $F=1$ leads to a relaxation rate of approximately 5/[s Amagat].
The general conclusion is that the pressure of O2 in a room-temperature cell must be below 6 mbar to achieve a spin lifetime of 200 s. In the case of Fig. 1, to achieve a lifetime of 3600 s requires a pressure of 0.3 mbar. This implies that the purity of the gas used to fill the cell to 233 mbar must have been about 99.9%, which is the typical level of commercial gases of non-UHP grade.
@@
A direct measure of the H relaxation rate would require the production of atomic H and a means to measure its pressure. Based on the foregoing, a total H2 pressure less than a few mbar appears as desirable in a cell, particularly given the destructive and autocatalytic behavior of atomic H.
@@ II Materials and Methods
To test the notions that have been discussed so far, the system shown in Fig. 6 was assembled in Fall 1991 at the University of Washington (these data are being reported for the first time). The system consisted of a selection of gases (99 + % purity) that could be introduced into an optical pumping cell contained in a three-layer magnetic shield, together with vacuum pumps and optical pumping components. The cell itself was of volume 750 $\ell$ (we refer to these as “1 liter cells") based on a design used to evaluate the optical pumping of 199Hg in larger volumes than had ever been done before, and was a key R&D component toward the development of the system described in [14].
The cell was coated with a generic pure PTFE lubricating spray in a hydrocarbon carrier. Cells with a coating made using a pure PTFE lubricating spray in a primarily hydro/fluorocarbon carrier (with minor components of isopropyl alcohol and acetone, without water) had better stability than other fluorinated coatings or any water-based Teflon coatings. An example is Elmer’s Slide-All® (which is no longer manufactured); although it was manufactured in a fast-evaporating water-free carrier, it contained a more complex polymer that has hydrogenous functional groups on the ends of PTFE polymer chains, and these groups are subject to attack by excited-state Hg or atomic H. A lubricating spray that uses the same molecule is still available, MicroCare® VDX Dry Lubricant Spray, Aerosol.
@@
where $R$ is the inner radius, and $h$ is the height. These coatings typically gave about 20 s of lifetime per cm mfp, which implies a lifetime of about 130 s for this large cell. The 1 mm fill hole, if complete depolarization occurs for 199Hg atoms entering that region, will lead to a relaxation rate of
@@
where $\bar{v}$ is the average Hg velocity and $A_{h}$ is the hole area, as based on the kinetic theory of atomic effusion. The observed cell lifetime was about 136 s, implying that the 199Hg atoms had a good chance of returning to the main volume with their polarization. This is because the capillary inside is coated with PTFE and which, given the smoothness of the surface, indicates that Hg atoms stick briefly to the surface and re-emitted in a random direction.
The operation of the system was straightforward. Natural Hg was supplied to the cell from a droplet until the transmitted light was reduced by a factor of 2. Various gases could be admitted and their pressures measured using a Hg manometer or McLeod gauge. A magnetic field of about 10 mG was applied along the light direction to polarize the atoms, and could be suddenly switched perpendicular to the light to observe the modulated light transmission due to the precessing atoms, as measured with a solar-blind photomultiplier. The magnetic shields had a substantial internal magnetic gradient; these shields were constructed from scrap materials and used for testing. The gradient limited the $T_{2}$ lifetime with gases at higher pressures; however, in those cases the longitudinal polarization pump-up and decay could be observed, giving $T_{1}\approx\tau_{0}$, the wall relaxation time.
@@
The main goal of the work was to assess the effect of hydrogen on the wall coating used in [14] which degraded with the application of high voltage, the application of which, we surmised, created atomic H due to inevitable microdischarges on dielectric surfaces. Previous work (see Fig. 1) suggested that oxygen might restore the cell relaxation lifetime. The secondary goal was to test for this possibility and determine how much oxygen could be used before limiting the 199Hg lifetime.
The tests were performed with PTFE instead of deuterated polystyrene (dps) [15] in part because in 1991 the final principal edm cell coating material used in [14] had not yet been decided. Typically, dps provides lifetimes of about 15 s per cm mfp, about as good as PTFE. Resources were not available to continue the 1991 studies with dps.
@@ III Results
Test 6 shows that hydrogen gas alone does not damage the wall coating. Tests 7-12 show that Hg in the presence of resonance light immediately degrades the cell, and removal of the gas does not repair the degraded lifetime. Given its stability, the degradation is probably not due to the atomic H being adsorbed on the surface but is more likely due to unterminated bonds on the wall coating. Atomic H has enough energy to do this damage, as shown in Table 1, and reported in [37, 36].
@@
Tests 15 and 17 indicate that the damage might be local as the buffer gas can tend to hide damage. The 254 nm light entered at exited the cell in a beam of about 2 cm diamter, so it is conceivable that the most significant coating damage was in these areas.
Table 2: Tests conducted with the apparatus shown in Fig. 6. All times have an uncertainty of 8 s unless noted otherwise.
@@
1 only Hg vapor 136 Cell starting lifetime
@@
15 increase to 110 torr " 145±$\pm 30$ pump up method so low signal; improvement
with buffer gas implies localized damage
@@
21 add 0.65 torr air $\approx 7$ relaxation rate $\approx 1100$/[s Amagat]
22 evacuate, add Hg 134 sat. Hg density, then evacuate to 1/2 abs.
23 evacuated, add Hg 110 evacuation was overnight $2\times 10^{-7}$ torr
24 1 hr irradiation 167 lifetime improving on irradiation
@@ IV Discussion and Conclusions
Based on the literature on photosensitized Hg reactions, together with our own observations, the results clearly indicate that hydrogen, through atomic H, is nearly certainly responsible for the degradation of 199Hg optically pumped magnetometry cells. Atomic H arises from the reaction of H2 with excited state Hg.
@@
Our 1991 measurement of the relaxation of 199Hg by oxygen gas agrees well with the most recent measurement reported in [48]. Given the salubrious effect of oxygen on cell lifetimes, the inclusion of a small amount will be beneficial; however, it will be consumed to form substances that stick to the wall or mix with wax. In addition, Hg will be consumed, so a source of 199Hg is needed within the cell, or ozone production must be suppressed. Given that of order $10^{13}$ photon per second are incident on a cell and that the photosensitize reaction in some cases has a quantum efficiency near unity, we might expect the same number of atoms or molecules lost per second. This corresponds to $10^{18}$ reaction per day, or 10 mbar of gas produced per day in a 5 $\ell$ cell.
@@
Hydrogen likely does its damage through an autocatalytic reaction, in that if there is any hydrogen, it will lead to photosensitized production of more hydrogen by breaking the bonds of hydrogenous materials on the cell walls. However, fluorinated waxes appear to not work as well as hydrocarbon waxes, but more studies are needed.
@@
Similarly, the seemingly anonymously long sticking times for 129Xe as reported in [40] can be explained by such states together with a low probability of sticking per wall collision, which is expected for Xe compared to Hg. Using microscopic reversibility and the calculable value of $\epsilon$ as the fraction of atoms stuck in the wall potential at any given time, if the dwell time can be calculated, perhaps using the Boltzmann transport equation for motion among the myriad of surface states, then the sticking probability is determined. Oddly enough, the reduced interaction of Xe probably means it stays on the surface longer, as the heating interactions that will lead to its release are likely smaller than those for Hg.
Regardless, there is another time scale that is better suited for the case of heavy atoms with many bound states, and that is the thermal wavelength divided by the average velocity, which of course is proportional $\hbar$ divided atom kinetic energy, which instead of the thermal wavelength, is the inverse of the thermal frequency.
An important consequence of the large number of relatively deep surface states (compared to $k_{b}T=0.025$ eV) is in regard to the assumption that a precise volume magnetic field average is provided by 199Hg co-magnetometry. For an extremely precise magnetometry system, as has been proposed, the surface dwell time (during which a substantial spin polarization survives) will need to be more carefully considered, as the dwell time can lead to corrections. For example, localized systematically generated magnetic fields are sensitive to the volume averaging. The gravitation offset of ultracold neutrons in neutron EDM experiments is perhaps a more serious problem; however, all such effects will eventually need consideration.
Based on the foregoing, to improve the durability of closed or open 199Hg optical pumping and detection cells (i.e., any similar cells used for fundamental experiments), some guidelines are as follows:
@@
    Avoid hydrogen gas either as an impurity in the CO buffer gas (oxygen can be removed, but water as an impurity will produce hydrogen which will pass through the amalgam) or generated by photosensitized reactions. Sealing of filled fused silica cells involves much heat and even UV light is produced, so care is needed in such operations;
@@
    For open cells used in the neutron EDM experiment that employs a 199Hg comagnetometer, the problem of lifetime degradation appears to be largely solved.[16] This is not surprising as the benzene ring is relatively immune from attack by atomic H [49], and exposure to ozone and oxygen lead to crosslinkages in the polymer.[50]
@@
    Photosensitized reactions can be suppressed by use of a quenching gas; CO has proven effective but there are better gases;
@@
    Fluorinated waxes or silanes with less hydrogen, e.g., Trichloro(3,3,3-trifluoropropyl)silane (CAS 592-09-6). A fully fluorinated silane would be of interest; however, none appear to be commercially available.
@@
    Deuterated polystyrene[15] remains one of the few options for UCN 199Hg comagnetometer experiments.

### v2

@@ I.5.1 Hydrogen surface states
Again, using the WKB approximation, the surface states of an H atom can be determined. The van der Waals coefficient is 2/3 that of Hg because the principal contributing optical transition for H is at a shorter wavelength. Also, the hard-sphere radius of H is smaller than that of Hg, so we can take $d_{0}\approx 0.12$ nm. Together with the smaller mass of H, this leads to the result that H has only four bound states at 0.08, 0.03, 0.01, and 0.003 $e$V. Although the energies are smaller than for Hg, $\epsilon$ is enhanced by a relatively large $\lambda_{T}$ compared to Hg. The net result for $T=293K$, is $\epsilon=2\times 10^{-6}$, about an order of magnitude smaller than for Hg. Nevertheless, this result implies a microscopically long residence time for atomic H on the surface, during which Hg atoms can interact with the H electronic magnetic moment.
We can estimate the effective pressure of H in the bound surface states, assuming that a fraction $\epsilon$ of the total H atoms are imprisoned in a region approximately $\delta=$ 0.3 nm thick. For a cell of radius $R$, the ratio of this surface volume to the cell volume is $3\delta/R$, so the relative densities are $\epsilon R/3\delta\approx 22$. For a given pressure $P$ of H gas in a 1 cm cell, the number density in the thin region near the cell wall corresponds to a pressure of about $22P$. Alternatively, with a pressure $P_{0}$ in cell volume, the pressure near the surface (similar to the isothermal atmosphere problem) should be $P_{0}e^{E_{max}/k_{b}T}=P_{s}$. For the maximum binding energy of 0.08 $e$ V for H from the WKB approximation, $e^{.08/.025}=24.5$ for $T=293$ K, which is in accord with the volume ratio determination.
Spin-polarization relaxation on the wall can be due to an enhanced interaction with H atoms or with the open bonds created when a surface is attacked by atomic H or excited Hg atoms. These open (covalent) bonds carry roughly an electron magnetic moment. The magnetic interaction that leads to relaxation is presented in [47] and can be easily modified for other magnetic moments.
@@ I.6 Relaxation in the Gas phase
Paramagnetic atoms and molecules, which include oxygen and free radicals, can exist in the buffer gas as impurities or the result of a 254 nm radiation induced photochemical reaction. Reactions with an excited Hg atom can release atomic H (an unpaired spin, which is technically a free radical) that exists in the states $F=0$ and $F=1$, in proportion to 1/4 and 3/4 of total H and $F+1$ presents an effective 1/2 of an electron magnetic moment due to the addition of the proton nuclear angular momentum.
The relaxation of nuclear spin-polarized $I=1/2$ 1 S0 atoms by gaseous paramagnetic atoms is theoretically analyzed by Jameson, Jameson, and Hwang [58] for the specific case of O2 and 129Xe, and further developed in [59], for O2 on both 3He and 129Xe. Rewritten here in terms of SI units and more generally for a colliding molecule (or atom) $X$ with total angular momentum $F$ representing the sum of the internal angular momentum of the constituent nuclei and the electron spins of $X$,
@@
$\gamma_{I}$ and $\gamma_{e}$ are the gyromagnetic ratios for the $I=1/2$ nucleus, and for the bare electron respectively; $g_{F}$ is the Landè $g-$ factor for $X$, and $\mu_{0}=4\pi\times 10^{-7}$. $d$ is the characteristic intermolecular interaction length, $\mu$ is the reduced mass of the colliding species, $k_{B}$ is the Boltzmann constant and $T$ is the absolute temperature (the factor in parentheses is the reciprocal of the mean relative velocity). $n_{0}$ is Loshmidt constant ($n_{0}=2.69\times 10^{25}$/m3 corresponding to the STP number density of an ideal gas), and $[X]$ is the density of the gas $X$ in Amagats. The factor $F(U,T)$ is determined by the interaction potential $U$ between the colliding species, and is unity at high temperature in the hard-core limit. In general, it is determined by integrating solutions of the radial Schrödinger equation with the potential $U$, together with an average over the Boltzmann momentum distribution. In [59] it is noted that the hyperfine Fermi contact term, proportional to $|\psi(0)|^{4}$, should generally be included in $F(U,T)$, where $\psi(0)$ is the amplitude of the electron wave function of the colliding molecule at the nucleus $I$. The interaction time (correlation time) for gas phase molecules is very short, so effects due to an applied magnetic field become small when the electron precession over the collision time results in a small directional change in the electron(s) spin angle(s); therefore, in all cases under consideration such effects can be ignored for the gas phase.
@@
In [59], the experimental relaxation rate for 3He in O2 is reported to be approximately $0.5$ /(s Amagat), which implies $F(U,T)=1.7$. For 129Xe, $F(U,T)\approx 6$ is estimated in [58], which implies a relaxation rate of 0.7/(s Amagat). (An Amagat is the number density of an ideal gas at a pressure of 1 bar or 760 torr at 273.16 K.)
@@
Data in [60] indicate that the spin polarization relaxation rate of 199Hg in O2 is 800/(s Amagat), which is approximately 1000 times faster than for the inert gases 3 He and 129 Xe. Using the analysis for inert gases implies that $F(U,T)=1.5\times 10^{4}$ for 199Hg . Because the two valence electrons of 199Hg enhance a hyperfine interaction between the atom and the molecule, the analysis for inert gases is not entirely accurate for Hg. Another possibility is that van der Waals molecules between 199Hg and O2 can be formed, in which case the needed three-body collisions would make the relaxation rate pressure dependence more complicated. This would be interesting to study.
The general conclusion is that the pressure of O2 in a room temperature cell must be less than the 6 $\mu$bar level to achieve $\tau_{w}>200$ s. This implies that the oxygen impurity of the nitrogen gas used to fill the cell in Fig. 1 was below 10 parts per million, consistent with its spectroscopic purity grade. Even with this high purity, the oxygen concentration in cells made using the same nitrogen had a rapid loss of Hg vapor under illumination with resonance radiation (the Hg filling pressure was 1.5 $\mu$ bars corresponding to the 0 vapor pressure, much lower than the oxygen pressure), as illustrated in Fig. 2.
@@
We can apply the foregoing analysis to the spin polarization relaxation of 199Hg by atomic H. Because H has a nuclear spin of 1/2, 3/4 of the atoms will be at hyperfine level with $F=1$ and $\gamma_{s}$ is reduced by a factor of two compared to O2. Comparing $F(U,T)$ with previous cases of inert gases suggests that $F(U,T)$ scales as the square root of the reduced mass of the 199Hg -H system, as expected from the Schrödinger equation. Scaling by $\mu$ and $\gamma_{s}^{2}$, and by the fraction of H atoms in $F=1$ leads to a relaxation rate for 199Hg nuclear spin polarization of approximately 5/(s Amagat).
@@
A direct measure of the H-induced spin polarization relaxation rate would require the production of atomic H and a means to measure its pressure. In light of such measurements, based on the foregoing, a total H2 pressure well below a few millibars appears to be necessary to stabilize $\tau_{w}$ in a cell, particularly given the destructive and autocatalytic behavior of atomic H.
@@ II Materials and Methods
To test the notions that have been discussed so far, the system shown in Fig. 4 was assembled in Fall 1991 at the University of Washington (these data are being reported for the first time). The system consisted of an optical pumping cell in a three-layer magnetic shield and optical components. The cell was connected to a vacuum system through a manifold that provided a selection of high purity gases (99.99% or better) or air that could be introduced into the cell. The cell itself was of volume 750 $\ell$ (we refer to these as “1 liter cells") based on a design used to evaluate the optical pumping of 199Hg in larger volumes than had ever been done before, and was a key R&D component toward the development of the system described in [20].
The cell was coated with a generic pure PTFE lubricating spray in a hydrocarbon carrier. Cells with a coating made using a pure PTFE lubricating spray in a primarily hydro/fluorocarbon carrier (with minor components of isopropyl alcohol and acetone, without water) had better stability than other partially fluorinated coatings or any water-based Teflon coatings. An example is Elmer’s Slide-All® (which is no longer manufactured); although it was manufactured with a fast-evaporating water-free carrier, it contained a more complex polymer with hydrogen-containing functional groups at the ends of the PTFE polymer chain, as shown in Fig. 5. These functional groups are subject to attack by excited-state Hg or atomic H and showed less stability than pure PTFE. A lubricating spray that uses the same molecule is still available, MicroCare® VDX Dry Lubricant Spray, Aerosol.
@@
where $R$ is the inner radius and $h$ is the height. These coatings typically gave about 20 s of $\tau_{w}$ per cm mfp, which implies a $\tau_{w}$ of about 130 s for this large cell. The 1 mm fill hole, if complete depolarization occurs for 199Hg atoms entering that region, will lead to a relaxation rate of
@@
based on the atomic flux determined by kinetic theory. The observed cell $\tau_{w}$ was about 136 s, implying that the 199Hg atoms had a good chance of returning to the main volume with their polarization. This is because the capillary inside is coated with PTFE and which, given the smoothness of the surface, indicates that Hg atoms stick briefly to the surface and re-emitted in a random direction.
The operation of the system was straightforward. Natural Hg was supplied to the cell from a droplet until the transmitted light was reduced by a factor of 2. Various gases could be admitted, and their pressures could be measured using a Hg manometer or McLeod gauge. A magnetic field of about 10 mG was applied along the light direction to polarize the atoms, and could be suddenly switched perpendicular to the light to observe the modulated light transmission due to the precessing atoms, as measured with a solar-blind photomultiplier. The magnetic shields had a substantial internal magnetic gradient; these shields were constructed from scrap materials and used for testing. The gradient limited $T_{2}$ $\tau_{w}$ with gases at higher pressures; however, in those cases, the longitudinal polarization pump-up and decay could be observed, giving $T_{1}\approx\tau_{0}$, the wall relaxation time.
@@
The main goal of the work was to assess the effect of hydrogen on the wall coating used in [20] for which $\tau_{w}$ degraded with the application of high voltage (HV). We surmised that the application of HV created atomic H as a result of inevitable microdischarges on the dielectric surfaces or dirt in the system. Previous work (see Fig. 1) suggested that oxygen might restore the cell $\tau_{w}$. The secondary goal was to test for this possibility and determine how much oxygen could be used before limiting the 199Hg spin polarization relaxation time.
The tests were performed with PTFE instead of deuterated polystyrene (DPS) [21] in part because in 1991 the final principal edm cell coating material used in [20] had not yet been decided. Typically, DPS provides $\tau_{w}$ of about 15 s per cm mfp, about as good as PTFE. Resources were not available to continue the 1991 studies with DPS.
@@ III Results
Test 6 shows that hydrogen gas alone does not damage the wall coating. Tests 7-12 show that Hg in the presence of resonance light immediately degrades the cell, and removal of the gas does not repair the degraded $\tau_{w}$. Given its stability, the degradation is probably not due entirely to the fact that atomic H is adsorbed on the surface but also to unterminated bonds on the wall coating. Atomic H has enough energy to do this damage, as shown in Table 1, and reported in [46, 45].
@@
Tests 15 and 17 indicate that the damage might be local as the buffer gas can tend to hide damage. The 254 nm light entered and exited the cell in a beam of about 2 cm diameter, so it is conceivable that the most significant coating damage was in these areas.
Table 2: Tests conducted with the apparatus shown in Fig. 4. All $\tau_{w}$ measurements have an uncertainty of 8 s unless noted otherwise.
@@
1 only Hg vapor 136 Cell starting $\tau_{w}$
@@
15 increase to 110 torr " 145±$\pm 30$ pump up method so low signal;
improvement with buffer gas
implies localized damage
@@
21 add 0.65 torr air $\approx 7$ O2 relaxation rate $\approx$840/(s Amagat)
22 evacuate, add Hg 134 fill to high Hg density,
then evacuate to 1/2 light absorption
23 evacuate, add Hg 110 evacuation overnight, to $2\times 10^{-7}$ torr
24 1 hr irradiation 167 $\tau_{w}$ improving on irradiation
@@ IV Discussion and Conclusions
Based on the literature on photosensitized Hg reactions, together with our own observations, the results clearly indicate that hydrogen, through atomic H, is nearly certainly responsible for much of the degradation of $\tau_{w}$ in 199Hg optically pumped magnetometry cells. Atomic H arises from the reaction of H2 with excited state Hg.
@@
Our 1991 measurement of the relaxation of 199Hg by oxygen gas agrees well with the most recent measurement reported in [60]. Given the salubrious effect of oxygen $\tau_{w}$, the inclusion of a small amount will be beneficial; however, it will be consumed to form substances that stick to the wall or mix with wall coating materials. In addition, Hg will be consumed to form HgO, so a source of 199Hg within the cell is needed, or ozone production must be suppressed. Given that of order $10^{13}$ photon per second are incident on a cell and that the photosensitized reaction in some cases has a quantum efficiency near unity, we might expect the same number of atoms or molecules lost per second. This corresponds to $10^{18}$ reaction per day, or 10 mbar of gas consumed per day in a 5 $\ell$ cell.
@@
Hydrogen likely does its damage through an autocatalytic reaction, in that if there is any hydrogen, it will lead to photosensitized production of more hydrogen by breaking the bonds of hydrogenous materials on the cell walls.
@@
Similarly, the seemingly anonymously long sticking times for 129Xe as reported in [50] can be explained through surface states together with a low probability of sticking per wall collision, which is expected for Xe compared to Hg. Using microscopic reversibility and the calculable value of $\epsilon$ as the average fraction of atoms stuck in the wall potential, the probability of sticking can be determined, which is of the order 1%.
Rather than the sticking time, there is another time scale that is better suited for the case of heavy atoms with many bound states. That time scale is the thermal wavelength divided by the average velocity, which is proportional $\hbar$ divided the kinetic energy of the atom. This time scale is the inverse of the “thermal frequency,” in analogy to the thermal wavelength.
The large number of relatively deep surface states (compared to $k_{b}T=0.025$ eV) brings into question the efficacy of 199Hg co-magnetometry. For an extremely precise magnetometry system, as has been proposed, the surface dwell time (during which a substantial spin polarization can survive) will need to be more carefully considered, as the dwell time can lead to corrections, e.g., localized systematically generated magnetic fields are sensitive to the volume averaging. The gravitation offset of ultracold neutrons in neutron EDM experiments is perhaps a more serious problem; however, all such effects will eventually need consideration.
Based on the foregoing, to improve the durability of closed or open 199Hg optical pumping and detection cells (i.e. any similar cells used for fundamental experiments), some guidelines are as follows:
@@
    Avoid hydrogen gas either as an impurity in the CO buffer gas (oxygen can be removed, but water as an impurity will produce hydrogen which will pass through the purification amalgam) or generated by photosensitized reactions. Sealing of filled fused silica cells involves much heat and even UV light is produced, so care is needed in such operations;
@@
    For open cells used in a neutron EDM experiment that employs a 199Hg comagnetometer, the problem of $\tau_{w}$ degradation appears to be largely solved, in particular by use of an oxygen gas rinse. [23, 22] The use of deuterated polystyrene offers some stability from attack by atomic H and other free radicals. This is in part due to the 1.6 eV energy required to disrupt the resonance stabilization of the phenyl ring. With this additional energy, an H atom cannot cause a C-H bond to break to form H-H; however, atomic H can be added to the phenyl ring, thus producing a cyclohexadienyl radical,[61] while exposure to ozone and oxygen leads to cross-linkages in the polymer.[62] This additional energy presents a barrier to direct reactions between benzene molecules and excited-state Hg atoms, as can be seen from the energies in Table 1. A study of the reaction between atomic O and benzene used the photosensitized reaction between Hg and N2O to produce O in situ in a reaction vessel, with the final amount of nitrogen as a measure of the total amount of reacted oxygen.[63]
    Pertaining to high voltage stability, the phenyl ring is attacked by positive ions that can bind to the electron cloud of the ring, leading to destabilization of the molecule due to disturbance of the resonance of the ring. This implies that DPS-coated cells require that high voltage can be applied without electrical discharges that produce positive ions. This also implies that DPS is less stable under high voltages than aliphatic compounds.
@@
    Photosensitized reactions in closed cells can be suppressed by use of a quenching gas; CO has proven effective but there are better gases;
@@
    Use fluorinated waxes or silanes with less hydrogen, e.g., Trichloro(3,3,3-trifluoropropyl)silane (CAS 592-09-6). A fully fluorinated silane would be of interest; however, none appears to be commercially available. Tests with fluorinated wax have not shown significant improvement but appear as promising.[43, 65]
@@
    Deuterated polystyrene (DPS) [21] remains one of the few options for UCN 199Hg comagnetometer experiments.